OpenStudy (mchilds15):

Create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. Equations: f(x)=2x+10 h(x)=3x+4 I chose x to equal 6 but how do I explain how the numbers won't repeat?

OpenStudy (anonymous):

f( h(x) ) is f(x), where x=h(x). 2(3x+4)+10 h( f(x)) is h(x) where x is f(x) 3(2x+10)+4

OpenStudy (anonymous):

2(3x+4)+10 = 6x+8+10 = 6x+18 3(2x+10)+4 = 6x+30+4 = 6x+34

OpenStudy (anonymous):

Using x=6 for both we get: 6(6)+18 = 54 6(6)+34 = 70 They will not result in the same number because the number following the 6x in both equations are different

OpenStudy (anonymous):

@Mchilds15

OpenStudy (mchilds15):

Oh thank you so much! I've been trying to figure this thing out for awhile... @larryboxaplenty